I know Surplus = Willingness to Pay - Explicit Cost, and I know that Opportunity Cost is Explicit Cost + Highest surplus (among remaining alternatives). Net Economic Benefit is Willingness to Pay - Opportunity Cost, however, I do not know how to put these all together to solve for the below problem. Any help would be appreciated.

(3) Consider the following decision problem, where is a real number and net economic benefits have already been calculated: AlternativeNet Economic Benefit Choice 1 Choice 2 3 Se lect the correct answer among the choices below.
Answer #1

Solution: 6

Working:

Net benefit = Total benefit - Total Cost

TB1 - TC = x

TB2 - TC =3

Thus, TC = 3 + TB2

TB1 - TC = x

TB1 + TB2 - 3 = x

Thus must be higher than 3